Question 7 Work out the following:

a) 3/4 - 7/9 + 2/3= 69/108 =


1
Expert's answer
2021-10-21T13:01:35-0400

"\frac{3}{4}-\frac{7}{9}+\frac{2}{3}=\frac{81}{108}-\frac{84}{108}+\frac{72}{108}=\frac{81-84+72}{108}=\frac{69}{108}=\frac{23}{36}"


Answer: "\frac{23}{36}"